# Subwavelength arrays of quantum emitters: nonlinearities enter the weak-drive regime, and lead to correlated subradiant states

### v3

@@ Appendix A Weakly-nonlinear theory of the weak-drive regime
The effective model is then that of a generalised parametric amplifier, with Hamiltonian $H=\delta(a^{\dagger}a+b^{\dagger}b)+\tilde{\delta}(a^{\dagger}b+b^{\dagger}a)+\lambda(a^{\dagger}a^{\dagger}+b^{\dagger}b^{\dagger}+\text{h.c.})+\tilde{\lambda}(a^{\dagger}b^{\dagger}+\text{h.c.})$, where we dropped the $k$ indices. This is typically encountered in the cases $\tilde{\delta},\tilde{\Lambda}=0$ and ${\delta},{\Lambda}=0$ [60], and is well known to undergo a parametric instability for a drive strength $|\lambda|^{2}>\delta^{2}+\Gamma^{2}$ ($|\tilde{\lambda}|^{2}>\tilde{\delta}^{2}+\Gamma^{2}$), accompanied by a divergence of steady-state populations. Assuming the same condition holds in our generalised case, and plugging $\lambda,\delta,\tilde{\lambda},\tilde{\delta}\sim\Omega^{2}$ and $\Gamma$ the parametric instability occurs for $\Omega\gtrsim N^{-\alpha/2}$ ($\Omega\gtrsim\exp(-N/2)$) for open (closed) boundary conditions. This is accompanied by a divergence of steady-state occupations.
We finally show that this condition holds even for $\tilde{\delta},\tilde{\Lambda},{\delta},{\Lambda}\neq 0$. We determine the instability in the Keldysh formalism, namely computing the Green’s function of the Keldysh-Nambu vector $\left(a_{+}^{\dagger}\;b_{+}^{\dagger}\;a_{+}\;b_{+}\;a_{-}^{\dagger}\;b_{-}^{\dagger}\;a_{-}\;b_{-}\right)$, where ++ and - are Keldysh indices. The inverse Green’s function, in frequency domain and for $\omega=0$, is given by
$G^{-1}(\omega=0)=\left(\begin{array}[]{cccccccc}-\delta-i\Gamma&-\tilde{\delta}&\lambda&\tilde{\lambda}&0&0&0&0\\ -\tilde{\delta}&-\delta-i\Gamma&\tilde{\lambda}&\lambda&0&0&0&0\\ \lambda^{*}&\tilde{\lambda}^{*}&-\delta-i\Gamma&-\tilde{\delta}&0&0&2i\Gamma&0\\ \tilde{\lambda}^{*}&\lambda^{*}&-\tilde{\delta}&-\delta-i\Gamma&0&0&0&2i\Gamma\\ 2i\Gamma&0&0&0&\delta-i\Gamma&\tilde{\delta}&-\lambda&-\tilde{\lambda}\\ 0&2i\Gamma&0&0&\tilde{\delta}&\delta-i\Gamma&-\tilde{\lambda}&-\lambda\\ 0&0&0&0&-\lambda^{*}&-\tilde{\lambda}^{*}&\delta-i\Gamma&\tilde{\delta}\\ 0&0&0&0&-\tilde{\lambda}^{*}&-\lambda^{*}&\tilde{\delta}&\delta-i\Gamma\\ \end{array}\right)$
Steady-state instabilities are then determined by the poles of $G$, which correspond to the zeros the determinant
$\operatorname{det}G^{-1}(\omega=0)=\left(\Gamma^{2}-|\lambda-\tilde{\lambda}|^{2}+(\delta-\tilde{\delta})^{2}\right)^{2}\left(\Gamma^{2}-|\lambda+\tilde{\lambda}|^{2}+(\delta+\tilde{\delta})^{2}\right)^{2}$
Inserting $\Gamma$ and $\lambda,\delta,\tilde{\lambda},\tilde{\delta}\sim\Omega^{2}$ in the determinant, both factors lead to the same instability condition as for the well-know cases above.
As a result of this analysis, we conclude that for $\Omega\gtrsim N^{-\alpha/2}$ ($\Omega\gtrsim\exp(-N/2)$) for open (closed) boundary conditions – and considering the most subradiant modes –, such perturbative approach is inadequate, and the steady-state is non-perturbative with respect to the non-interacting (classical) prediction. The exponent $\alpha$ was found to be $\alpha\approx 3$ in 1D, and $\alpha\approx 6$ or $3$ in 2D [25].
Appendix B Nonequilibrium steady-state DMFT
@@
In the present case, the effective single-site model obtained in DMFT (often called impurity model) is a generalized spin-boson model, which in the steady state is described by the time-translation invariant Keldysh action
@@
Appendix C Schemes for solving the DMFT equations
@@
where $\alpha\in[0,1]$ is a mixing parameter. This method allowed to reach smaller drive strengths, yet requiring an increasing smaller $\alpha$ the smaller the drive strength, and could not converge below a certain point. Indeed, a linear mixing approach still does not guarantee convergence to all solutions [63].
@@
An example of convergence at weak-drives is shown in Fig. S.1, where the DMFT error at the $m$-th iteration is defined as
@@
Figure S.1: DMFT error as a function of number of DMFT iterations, starting close to the previously found solution at $k_{0}=2$ and $\Omega=0.3,\Delta=0$. Here $dt=0.1$ and $t_{\rm max}=200$.
@@
Appendix D Exact limits of DMFT
Here we discuss two exact limits of DMFT, in the context of our problem. The first is the usual atomic limit. In this case this corresponds to independent emitters, in which the impurity problem becomes equivalent to a single site of the original lattice problem: the impurity action (10) reduces to a single-site action with $b=0$ and $\mathcal{W}=1/N\sum_{k}W_{k}$ describing the local dissipation. We note that in this limit the NCA solver used here [51, 52, 61] exactly reduces to a Lindblad master equation, introducing no further approximations on the solution of the single-site problem.
In addition, the DMFT becomes exact in the low-density regime in which the model reduces to non-interacting bosons discussed in the main text. To show this, consider its derivation reported in [51], the DMFT is here formulated on a lattice problem equivalent to Eq. (1), in which auxiliary bosonic degrees of freedom are introduced by decoupling the non-local terms (the $V_{ij}$ and $\Gamma_{ij}$ terms) with a Hubbard-Stratonovich transformation. The resulting boson-atom lattice problem action reads $S_{\mathrm{HS}}=S_{0}+S_{\phi\phi}+S_{\phi\sigma}$, where $S_{0}=\sum_{r}S_{0,r}$ is the action of decoupled emitters at position $r$, $S_{\phi\phi}$ is a free-bosons action in terms of a Nambu vector of complex bosonic fields $\phi_{r}$
@@ Appendix D Exact limits of DMFT
and $W_{rr^{\prime}}(t-t^{\prime})$ is the Fourier transform of (15). $S_{\phi\sigma}$ describes a local linear coupling term between bosons and atoms (with spin Nambu fields $\varsigma$ defined as in (10))
@@
In the case in which non-linearities can be discarded, the spin variables can be replaced by bosonic ones, resulting into a gaussian action of a single atom that can be integrated out exactly. Importantly, this results into a self-energy for the auxiliary bosons that is local in space $\Pi_{rr^{\prime}}=\Pi_{\rm loc}\delta_{rr^{\prime}}$– this is the case because the atoms are only coupled to each other through the auxiliary bosons. One can also show that the boson-atom lattice model, being entirely quadratic, can be exactly mapped onto the impurity model (10) with a cavity construction where all other sites are integrated out. Since these are the two assumptions on which DMFT rests [70, 71], this approach becomes exact in this non-interacting regime. We confirmed this numerically, as shown by Fig. S.2 where the local Green’s functions computed via DMFT and the non-interacting bosons model agree, outside of the subwavelength regime (for $k_{0}a>\pi$). The two instead deviate in the subwavelength regime ($k_{0}a<\pi$), as it is later shown in Fig. S.3, because non-linearities become important there as pointed out in the main text. We remark that the non-crossing approximation (NCA) used to solve the impurity problem introduces a further approximation, which is controlled when the emitters are weakly coupled. This eventually limits our DMFT/NCA approach at low $k_{0}a$ values and weak drives, for which the emitters behave collectively.
@@
Figure S.2: The local atomic Keldysh Green’s function defined as in Fig. S.3 computed from a non-interacting theory for $\Omega\rightarrow 0$ (dashed line) and in DMFT (solid), for $k_{0}a=\pi$ and $\Delta=0$. This deviates from a Lorentzian shape for a single atom due the dipolar interactions between the emitters.
Appendix E Linear stability equations
@@
The homogeneous DMFT equation (14) can be generalized to allow a momentum and frequency dependent effective field, while keeping the effective bath homogeneous, allowing to study the linear stability of such perturbations.
@@
The field must then satisfy the self-consistent condition $b_{k}(\omega)=\left[\mathcal{W}^{R}(\omega)-W_{k}^{R}(\omega)\right]\langle{\varsigma_{k}\left(\omega\right)}\rangle$ and, assuming a perturbation around a homogeneous value $b_{k}(\omega)=b\delta(k)\delta(\omega)+\delta b_{k}(\omega)$, the steady-state expectation $\langle{\varsigma_{k}(\omega)}\rangle=\langle{\varsigma}\rangle\delta(k)\delta(\omega)+\langle{\delta\varsigma_{k}(\omega)}\rangle$ can be obtained from linear response $\langle{\delta\varsigma_{k}(\omega)}\rangle=-\chi_{k}^{R}(\omega)\delta b_{k}(\omega)$. Here “RR” indicates the retarded components of $\mathcal{W}$, representing the effective environment, $\chi$ is the local atomic Green’s function and $W_{k}$ a matrix of the coherent and dissipative couplings (4) and (5) defined in Eq. (15). Then a finite $\delta b_{k}(\omega)$ can form if the following condition is satisfied:
@@
Appendix F Relaxation of local correlation functions in the non-interacting bosons approximation
In the simplest non-interacting bosons approximation, the steady-state single-particle Green’s functions are easily obtained in a Keldysh formalism. Assuming the thermodynamic limit $N\rightarrow\infty$, in the limit of vanishing drive $\Omega\rightarrow 0$ the problem becomes diagonal in Nambu space and one gets the inverse Green’s function in the basis of classical and quantum fields
@@ Appendix F Relaxation of local correlation functions in the non-interacting bosons approximation
Even in a non-interacting theory, where the decay rate of non-radiative modes is zero, such correlation function could in principle still relax to a steady state by dephasing of the different modes, constituting a continuum. In Fig. S.3 we show that nevertheless these fail to relax. An interesting question then is whether taking non-linearities into account, which can provide an effective lifetime for the non-radiative modes, allows these quantities to relax.
In DMFT/NCA we find that these quantities do relax and correspond to a smoothened version in frequency of those predicted by a non-interacting theory, as shown in Fig. S.3. These predictions might be nevertheless limited by inaccuracies of the non-crossing approximation in this regime. The prediction of the non-interacting theory where an artificial damping is added was also useful in the present work to initialize the DMFT.
Figure S.3: The local atomic Keldysh Green’s function ${\rm i}\chi_{11}^{K}(\tau)=\lim_{t\rightarrow\infty}\langle{\left\{\sigma^{-}(t+\tau),\sigma^{+}(t)\right\}}\rangle-\left\{\langle{\sigma^{-}(t+\tau)}\rangle,\langle{\sigma^{+}(t)}\rangle\right\}$ computed from a non-interacting theory for $\Omega\rightarrow 0$ and in DMFT in real time (a) and frequency (b), for the same parameters as in Fig. 2 (c).

### v4

@@ Abstract
The noninteracting part of the effective model is then that of a two-modes parametric amplifier [58], with Hamiltonian $H=\delta(a^{\dagger}a+b^{\dagger}b)+{\lambda}(a^{\dagger}b^{\dagger}+\text{h.c.})$ – where we dropped the $k$ indices. This is well known [58] to undergo a parametric instability for a drive strength $|\lambda|^{2}>\delta^{2}+\Gamma^{2}$, accompanied by a divergence of steady-state moments (its populations and anomalous correlators). In our case, this instability occurs for $\Omega\gtrsim N^{-\alpha/2}$ ($\Omega\gtrsim\exp(-N/2)$) for open (closed) boundary conditions. For the most subradiant modes, the exponent $\alpha$ was found to be $\alpha\approx 3$ in 1D, and $\alpha\approx 6$ or $3$ in 2D [4].
Beyond this threshold, the noninteracting model is dynamically unstable, and it therefore cannot be used as the basis of a perturbative analysis. The steady-state must then be obtained using non-perturbative approaches.
SM Note B Nonequilibrium steady-state DMFT
@@
In the present case, the effective single-site model obtained in DMFT (impurity model) is a generalized spin-boson model, which in the steady state is described by the time-translation invariant Keldysh action
@@
SM Note C Schemes for solving the DMFT equations
@@
where $\alpha\in[0,1]$ is a mixing parameter. This method allowed to reach smaller drive strengths, yet requiring an increasingly-smaller $\alpha$ the smaller the drive strength, and could not converge below a certain point. Indeed, a linear mixing approach still does not guarantee convergence to all solutions [71].
@@
An example of convergence at weak-drives is shown in Fig. S1, where the DMFT error at the $m$-th iteration is defined as
@@
Figure S1: DMFT error as a function of number of DMFT iterations, starting close to the previously found solution at $k_{0}=2$ and $\Omega=0.3,\Delta=0$. Here $dt=0.1$ and $t_{\rm max}=200$.
@@
SM Note D Exact limits of DMFT
Here we discuss two exact limits of DMFT, in the context of our formulation. The first is the usual atomic limit – here the limit of independent emitters. In this limit the impurity problem becomes equivalent to a single site of the original lattice problem: the impurity action (10) reduces to a single-site action with $b=0$ and $\mathcal{W}=1/N\sum_{k}W_{k}$ describing the local dissipation. In addition, in this limit the NCA solver we use [55, 54, 57] exactly reduces to a Lindblad master equation, introducing no further approximations on the solution of the single-site problem.
In addition, DMFT becomes exact at low-densities, where the model reduces to a model of noninteracting bosons. To show this, we refer to the derivation of the approach, reported in [55], or in [38]. DMFT is formulated on a lattice problem equivalent to Eq. (1), in which auxiliary bosonic degrees of freedom are introduced by decoupling the non-local terms (the $V_{ij}$ and $\Gamma_{ij}$ terms) with a Hubbard-Stratonovich transformation. The resulting boson-spin lattice problem action reads $S_{\mathrm{HS}}=S_{0}+S_{\phi\phi}+S_{\phi\sigma}$, where $S_{0}=\sum_{r}S_{0,r}$ is the action of decoupled spins at position $r$, $S_{\phi\phi}$ is a free-bosons action in terms of a Nambu vector of complex bosonic fields $\phi_{r}$
@@
and $W_{rr^{\prime}}(t-t^{\prime})$ is the Fourier transform of (15). $S_{\phi\sigma}$ describes a local linear coupling term between bosons and spins (with spin Nambu fields $\varsigma$ defined as in (10))
@@
In the case in which nonlinearities can be discarded, the spin variables can be replaced by bosonic ones, resulting into a gaussian action of a single atom that can be integrated out exactly. Importantly, this results into a self-energy for the auxiliary bosons that is local in space $\Pi_{rr^{\prime}}=\Pi_{\rm loc}\delta_{rr^{\prime}}$– this is the case because the atoms are only coupled to each other through the auxiliary bosons. One can also show that the boson-spin lattice model, being entirely quadratic, can be exactly mapped onto the impurity model (10) with a cavity construction where all other sites are integrated out. Since these are the two assumptions on which DMFT rests [22, 3], this approach becomes exact in this noninteracting regime. This is shown numerically in Fig. S2, where the local Green’s functions computed via DMFT and the noninteracting bosons model agree, outside of the subwavelength regime (for $k_{0}a>\pi$). We remark that the non-crossing approximation (NCA) of the impurity problem introduces a further approximation, but here this is controlled because the interactions between different emitters are relatively weak. The two quantities instead are shown to deviate in the subwavelength regime where interactions are stronger ($k_{0}a<\pi$) in Fig. S3. This eventually limits our DMFT/NCA approach at low $k_{0}a$ values and weak drives, for which the emitters behave strongly collectively.
@@
Figure S2: The local atomic Keldysh Green’s function defined as in Fig. S3 computed from a noninteracting theory for $\Omega\rightarrow 0$ (dashed line) and in DMFT (solid), for $k_{0}a=\pi$ and $\Delta=0$. This deviates from a Lorentzian shape for a single atom due the dipolar interactions between the emitters.
SM Note E Linear stability equations
@@
The homogeneous DMFT equation (14) can be generalized to allow a momentum and frequency dependent effective field, allowing to study the linear stability of such perturbations. Note that, in order to obtain a condition that can be evaluated from the homogeneous solution only, we don’t also allow for an inhomogeneous perturbation of the effective bath, that we keep instead homogeneous.
@@
The field must then satisfy the self-consistent condition $b_{k}(\omega)=\left[\mathcal{W}^{R}(\omega)-W_{k}^{R}(\omega)\right]\langle{\varsigma_{k}\left(\omega\right)}\rangle$ and, assuming a perturbation around a homogeneous value $b_{k}(\omega)=b\delta(k)\delta(\omega)+\delta b_{k}(\omega)$, the steady-state expectation $\langle{\varsigma_{k}(\omega)}\rangle=\langle{\varsigma}\rangle\delta(k)\delta(\omega)+\langle{\delta\varsigma_{k}(\omega)}\rangle$ can be obtained from linear response $\langle{\delta\varsigma_{k}(\omega)}\rangle=-\chi_{k}^{R}(\omega)\delta b_{k}(\omega)$. Here “RR” indicates the retarded components of $\mathcal{W}$, representing the effective environment, $\chi$ is the local atomic Green’s function and $W_{k}$ a matrix of the coherent and dissipative couplings (4) and (5) defined in Eq. (15). Then, we obtain that a finite $\delta b_{k}(\omega)$ can form if the following condition is satisfied:
@@
SM Note F Relaxation of local correlation functions
In the simplest noninteracting bosons approximation, the steady-state single-particle Green’s functions are easily obtained in a Keldysh formalism. Assuming the thermodynamic limit $N\rightarrow\infty$, in the limit of vanishing drive $\Omega\rightarrow 0$ the problem becomes diagonal in Nambu space and one gets the inverse Green’s function in the basis of classical and quantum fields
@@
Even in a noninteracting theory, where the decay rate of non-radiative modes is zero, such correlation function could in principle still relax to a steady state by dephasing of the different modes, constituting a continuum. In Fig. S3 we show that nevertheless these fail to relax. An interesting question then is whether taking nonlinearities into account, which can provide an effective lifetime for the non-radiative modes, allows these quantities to relax.
In DMFT/NCA we find that these quantities do relax and correspond to a smoothened version in frequency of those predicted by a noninteracting theory, as shown in Fig. S3. These predictions might be nevertheless limited by inaccuracies of the non-crossing approximation in this regime. The prediction of the noninteracting theory where an artificial damping is added was also useful in the present work to initialize the DMFT.
Figure S3: The local atomic Keldysh Green’s function ${\rm i}\chi_{11}^{K}(\tau)=\lim_{t\rightarrow\infty}\langle{\left\{\sigma^{-}(t+\tau),\sigma^{+}(t)\right\}}\rangle-\left\{\langle{\sigma^{-}(t+\tau)}\rangle,\langle{\sigma^{+}(t)}\rangle\right\}$ computed from a noninteracting theory for $\Omega\rightarrow 0$ and in DMFT in real time (a) and frequency (b), for the same parameters as in Fig. 2 (c).